A gemstone has a mass of 6.230 g. Placed in a graduated cylinder, the volume increases from 45.0 mL to 46.5 mL. What is the volume of the gemstone and what is its density?
1
Expert's answer
2020-09-23T04:43:03-0400

Ans.

The volume of gemstone = increased volume in graduated cylinder = (46.5-45) ml = 1.5 ml

Density of gemstone = (6.230/1.5)g/ml = 4.153 g/ml (Ans.)